# Soft-QMIX: Integrating Maximum Entropy For Monotonic Value Function Factorization

### v2

@@ 5.4 Ablation Study
We evaluate the effectiveness of different components in our algorithm through an ablation study. We begin with QMIX, which utilizes epsilon-greedy exploration. We then introduce +entropy, which incorporates softmax sampling (employing $\text{softmax}(Q_{i})$ as the policy) along with a maximum entropy objective function to the QMIX framework. Note that this version of the algorithm does not ensure convergence, leading to a performance decline in two out of three scenarios. The subsequent modification, +function f, integrates the function $f_{i}$ (using $\text{softmax}(f_{i}(Q_{i})$) as the policy) building upon the +entropy setup. Empirical results show that function f yields the most substantial performance improvement across all components. Our study culminates with Soft-QMIX, our finalized algorithm. For each modification, we conduct four experiments across three seeds, reporting the average winning rate and its standard deviation in our results table.
Table 4: The ablation studies reveal that each component enhances performance, with function $f$ contributing most significantly to the performance improvement.
eval win rate QMIX +entropy +function f Soft-QMIX(ours)
protoss 5vs5 0.68 $\pm$ 0.04 0.73 $\pm$ 0.05 0.74 $\pm$ 0.04 0.76 $\pm$ 0.04
terran 5vs5 0.68 $\pm$ 0.05 0.67 $\pm$ 0.04 0.70 $\pm$ 0.05 0.72 $\pm$ 0.04
zerg 5vs5 0.41 $\pm$ 0.05 0.40 $\pm$ 0.04 0.48 $\pm$ 0.05 0.51 $\pm$ 0.04
@@ 6 Conclusion
In this paper, we introduced Soft-QMIX, an innovative approach within the CTDE framework, integrating the maximum entropy RL framework with the QMIX-based credit assignment mechanism. Theoretically, we established the monotonic improvement and convergence of Soft-QMIX to a globally optimal solution. Experimentally, our algorithm can solve matrix games and demonstrate state-of-the-art performance on the SMAC-v2 benchmark. Currently, Soft-QMIX is a value-based method, thus one of its limitations is that it can currently only be deployed in tasks with discrete action spaces. In the future, we will try to extend it to tasks with continuous action spaces.
@@ Acknowledgment
Appendix A Soft-QMIX
Algorithm 1 Soft-QMIX
1:  Input: $\theta,\phi$
2:  Initiate: $\bar{\theta}\leftarrow\theta,D\leftarrow\emptyset$
3:  for each iteration do
4:     for each environment step do
5:        get $\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t})$ via Eq. 15
6:        $u^{i}_{t}\sim\pi_{i}(u^{i}_{t}|o^{i}_{t},s_{t}),\forall i\in\{1,..,n\}$
7:        $s_{t+1}\sim p(s_{t+1}|s_{t},\mathbf{u}_{t})$
9:     end for
10:     for each gradient step do
11:        calculate $\mathcal{T}^{\pi_{jt}}_{\lambda}Q_{tot}$ via Eq. 24
12:        $\theta\leftarrow\theta-\lambda_{Q}\nabla_{\theta}L_{Q}(\theta)$
13:        $\phi\leftarrow\phi-\lambda_{\pi}\nabla_{\phi}L_{\pi}(\phi)$
14:        $\bar{\theta}\leftarrow\tau\theta+(1-\tau)\bar{\theta}$
15:     end for
16:  end for
Appendix B Individual-Global-Independence (IGI)
In this section, we first introduce two concepts, IGI and IGO, where IGI is proposed in our work, and IGO was proposed by A. We then explain that IGI is a sufficient condition for IGO.
Definition B.1.
Individual-Global-Independence (IGI) For joint policy $\pi_{jt}$, if their exist individual policies $\left[\pi_{i}\right]_{i=1}^{n}$ such that the following holds:
$\pi_{jt}(\mathbf{u}|s)=\prod_{i=1}^{n}\pi_{i}(u^{i}|o^{i}).$ (16)
Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGI for $\pi_{jt}$ under $s_{t}$.
Definition B.2.
Individual-Global-Optimal (IGO) For optimal joint policy $\pi^{*}_{jt}$, if their exist individual optimal policies $\left[\pi^{*}_{i}\right]_{i=1}^{n}$ such that the following holds:
$\pi^{*}_{jt}(\mathbf{u}|s)=\prod_{i=1}^{n}\pi^{*}_{i}(u^{i}|o^{i}).$ (17)
Then, we say that $[\pi_{i}]_{i=1}^{n}$ satisfies IGO for $\pi_{jt}$ under $s_{t}$.
IGO defines that the optimal joint policy equals the product of optimal local policies. In IGI, it is defined that any joint policy should equal the product of local policies. If finding the optimal policy is viewed as an optimization problem, IGI defines the feasible set of this problem. Hence, any policy not equaling the product of local policies is not a feasible solution and thus cannot be optimal. Therefore, all optimal policies will satisfy the joint policy equals the product of local optimal policies, thereby meeting the IGO condition.
Note that under the assumption of a decentralized execution setting, IGI and IGO are equivalent.
Appendix C Joint Policy Iteration
Lemma C.1 (Joint Soft Policy Evaluation).
Let $Q^{0}_{tot}:S\times\mathbf{U}\to\mathbb{R}$ be a bounded mapping where $|\mathbf{U}|<\infty$. Define the iterative process $Q^{k+1}_{tot}=\mathcal{T}^{\pi_{jt}}_{\lambda}(Q^{k}_{tot})$. It follows that the sequence $\{Q^{k}_{tot}\}_{k=0}^{\infty}$ converges to the joint soft-Q-function for policy $\pi_{jt}$ as $k\rightarrow\infty$.
@@ Appendix C Joint Policy Iteration
Initiate by defining the entropy-augmented reward function:
$r_{\pi_{jt}}(s_{t},\mathbf{u}_{t})\coloneqq r(s_{t},\mathbf{u}_{t})+\mathbb{E}_{s_{t+1}}\left[H[\pi_{jt}(\cdot|s_{t+1})]\right].$ (18)
Follow previous work [9], we choose a lower $\lambda$ and can assume that $\mathcal{T}^{\pi_{jt}}_{\lambda}$ is an unbiased estimator of $\mathcal{T}^{\pi_{jt}}$. Subsequently, the update mechanism is articulated as:
$Q_{tot}(s_{t},\mathbf{u}_{t})$ $\leftarrow r_{\pi_{jt}}(s_{t},\mathbf{u}_{t})+\gamma\mathbb{E}_{s_{t+1},\mathbf{u}_{t+1}\sim\pi_{jt}}[Q_{tot}(s_{t+1},\mathbf{u}_{t+1})].$ (19)
Conclude by invoking the established convergence theorems pertinent to policy evaluation as explicated by Sutton and Barto. ∎
Lemma C.2 (Joint Soft Policy Improvement).
@@
Let $\pi_{jt}^{\text{old}}\in\Pi$ and let $Q^{\pi_{jt}^{\text{old}}}$ and $V^{\pi_{jt}^{\text{old}}}$ be the corresponding soft state-action value and soft state value. The update rule of $\pi_{\text{new}}$ can be defined as
@@
$=\arg\min_{\pi^{\prime}\in\Pi}J_{\pi_{jt}^{\text{old}}}\left(\pi^{\prime}(\cdot\mid s_{t})\right).$
@@
where we can repeatedly expand $Q^{\pi_{jt}^{\text{old}}}$ on the RHS by applying the soft Bellman equation and the bound in Equation 23. Convergence to $Q^{\pi_{jt}^{\text{new}}}$ follows from Lemma B.1.
Theorem C.3 (Joint Soft Policy Iteration).
By continually applying joint soft policy evaluation and joint soft policy improvement to any $\pi_{jt}\in\Pi$ converges to a policy $\pi^{*}_{jt}$ such that $Q^{\pi^{*}_{jt}}(s_{t},\mathbf{u}_{t})\geq Q^{\pi_{jt}}(s_{t},\mathbf{u}_{t})$ for all $\pi_{jt}\in\Pi$ and $(s_{t},\mathbf{u}_{t})\in S\times\mathbf{U}$, assuming $|\mathbf{U}|<\infty$.

### v3

@@ 6.4 Ablation Study
Figure 5: Comparison of exploration strategies in CTDE MARL: ME-QMIX with maximum entropy vs. QMIX with epsilon-greedy and extended epsilon annealing period.
Benefits of entropy-driven exploration: To further demonstrate the benefits of using the maximum entropy framework for exploration, we compare ME-QMIX with a modified version of QMIX, which incorporates significantly higher exploration by applying an extended epsilon annealing period in the epsilon-greedy method, on Protoss 5v5. Figure 5 illustrates that simply increasing the epsilon annealing time for exploration does not improve QMIX’s performance. ME-QMIX achieves superior performance, with the systematic exploration enabled by the maximum entropy framework.
QMIX +Entropy +OPT
Protoss 5vs5 0.68 $\pm$ 0.04 0.73 $\pm$ 0.05 0.74 $\pm$ 0.04
Terran 5vs5 0.68 $\pm$ 0.05 0.67 $\pm$ 0.04 0.70 $\pm$ 0.05
Zerg 5vs5 0.41 $\pm$ 0.05 0.40 $\pm$ 0.04 0.48 $\pm$ 0.05
Table 3: Our algorithm (+OPT) vs. its two ablated versions: +Entropy, obtained by removing OPTs, and QMIX, obtained by further removing softmax local policies.
The necessity of using OPTs: In Table 3, we compare QMIX with its two variants: +Entropy and +OPT. Both variants learn local stochastic policies based on Q-functions trained with the QMIX objective. Specifically, +OPT corresponds to our proposed ME-IGM, while +Entropy removes the OPTs from ME-IGM (i.e., no OPTs in Equations (4), (9), and (10)). We test each algorithm on three SMAC-v2 scenarios across four random seeds, reporting the average win rate along with its standard deviation in the table. Without OPTs, +Entropy suffers from the misalignment problem discussed in Section 3 and fails to ensure monotonic policy improvement, as analyzed in Section 4.2, leading to a performance decline in two out of three scenarios. On the other hand, +OPT achieves a substantial performance improvement across all evaluated scenarios.
(a) Impact of the learning rate for $\alpha_{\omega}$
(b) Impact of the target entropy $\bar{\mathcal{H}}$
Figure 6: The impact of different hyperparameters for training $\alpha_{\omega}$ on the performance of ME-QMIX. (a) A lower learning rate for $\alpha_{\omega}$ corresponds to increased exploration. (b) ME-QMIX’s performance is not sensitive to the choice of the target entropy $\bar{\mathcal{H}}$ in Equation (10).
ME-QMIX (MLP) ME-QMIX (OPT)
Protoss 5v5 0.58±0.030.58$\pm$ 0.030.58 ± 0.03 0.74±0.040.74$\pm$ 0.040.74 ± 0.04
Terran 5v5 0.61±0.040.61$\pm$ 0.040.61 ± 0.04 0.70±0.050.70$\pm$ 0.050.70 ± 0.05
Zerg 5v5 0.41±0.040.41$\pm$ 0.040.41 ± 0.04 0.48±0.050.48$\pm$ 0.050.48 ± 0.05
Table 4: ME-QMIX (OPT) vs. its ablated version: ME-QMIX (MLP), where OPT is replaced with a 3-layer MLP.
To further validate the design of the order-preserving transformation, we replace the OPT layers in ME-QMIX with a 3-layer MLP, which does not enforce non-negative weights, referring to this variant as ME-QMIX (MLP). As shown in Table 4, our method, ME-QMIX (OPT), consistently outperforms ME-QMIX (MLP) across three different SMAC-v2 scenarios. This underscores the importance of the order-preserving transformation (i.e., Eq. (2)) in resolving the misalignment between local policies and the maximum global Q-value, thereby enhancing overall performance.
The impact of hyperparameters: Compared to QMIX, ME-QMIX introduces two additional hyperparameters: the learning rate and the target entropy $\bar{\mathcal{H}}$ for updating $\alpha_{\omega}$ in Eq. (10). We set the other hyperparameters of ME-QMIX following the design choices of QMIX. To analyze the impact of these additional parameters, we conduct a parameter scan on the learning rate of $\alpha_{\omega}$ and the target entropy $\bar{\mathcal{H}}$. Figure 6(a) illustrates that the learning rate of $\alpha_{\omega}$ influences the exploration level. A lower learning rate results in higher policy entropy at the beginning of training, promoting exploration. While this initially leads to weaker performance (e.g., the blue curve in Figure 6(a)), it eventually achieves higher returns. Furthermore, the results in Figure 6(b) suggest that the algorithm’s performance is not sensitive to the choice of the target entropy. Notably, in this study, we did not fine-tune configurations for each individual testing scenario. Please check the Appendix for the tables of hyperparameters.
@@ 7 Conclusion
In this paper, we introduce ME-IGM, a novel maximum entropy MARL approach that applies the CTDE framework and satisfies the IGM condition. We identify and address the misalignment problem between local policies and the maximum global Q-value (in maximum entropy MARL) by introducing an order-preserving transformation (OPT). We further propose a theoretically sound and straightforward objective for updating these transformation operators. Empirically, we demonstrate the effectiveness of ME-IGM on matrix games and its state-of-the-art performance on the challenging Overcooked and SMAC-v2 benchmarks. However, as a value-based method, ME-IGM is currently limited to tasks with discrete action spaces. In future work, we plan to extend its use to tasks involving continuous action spaces.
@@ Acknowledgment
[…]
@@
The one-step TD error can be expressed in the following form:
$\delta_{t}^{V}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})$ (13)
Next, the k-step TD error can be represented in the following form:
$\epsilon_{t}^{(1)}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma V_{tot}(s_{t+1})=\delta^{V}_{t}$ (14)
$\epsilon_{t}^{(2)}=-Q_{tot}(s_{t},\mathbf{u}_{t})+r_{t}+\gamma r_{t+1}{-\log\pi_{jt}(\mathbf{u_{t+1}}|s_{t+1})}+\gamma^{2}V_{tot}(s_{t+1})=\delta^{V}_{t}+\gamma\delta^{V}_{t+1}$
$\epsilon_{t}^{(k)}=\sum_{l=0}^{k-1}\gamma^{l}\delta^{V}_{t+l}$
$\epsilon_{t}^{td(\lambda)}=(1-\lambda)\left(\epsilon_{t}^{(1)}+\lambda\epsilon_{t}^{(2)}+\lambda^{2}\epsilon_{t}^{(3)}+\cdots\right)$ (15)
$=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{V}$
Appendix B B. Proof of Theorem 4.1
Lemma B.1 (Joint Soft Policy Improvement).
@@
Let $\pi_{jt}^{\text{old}}\in\Pi$ and let $Q^{\pi_{jt}^{\text{old}}}$ and $V^{\pi_{jt}^{\text{old}}}$ be the corresponding soft state-action value and soft state value. The update rule of $\pi_{\text{new}}$ can be defined as:
@@
$=\arg\min_{\pi^{\prime}\in\Pi}J_{\pi_{jt}^{\text{old}}}\left(\pi^{\prime}(\cdot\mid s_{t})\right),$
where $Z^{\pi_{jt}^{\text{old}}}(s_{t})$ is the normalization term.
@@
where we can repeatedly expand $Q^{\pi_{jt}^{\text{old}}}$ on the RHS by applying the soft Bellman equation, progressing from the second line to the final one.
Theorem B.2.
Denote the policy before the policy improvement as $\pi_{\text{old}}$ and the policy achieving the optimality of the improvement step as $\pi_{\text{new}}$. Updating the policy with Equation (4) ensures that $Q^{\pi_{\text{old}}}_{tot}(s_{t},\mathbf{u_{t}})-\epsilon\leq Q_{tot}^{\pi_{\text{new}}}(s_{t},\mathbf{u_{t}}),\forall s_{t},\mathbf{u_{t}}$. The gap $\epsilon$ satisfies:
$\epsilon\leq\frac{\gamma}{1-\gamma}\mathbb{E}_{s_{t+1}\sim\mathcal{P}}\big{[}C\sqrt{2D_{KL}\big{(}\pi_{\text{old}}(\cdot|s_{t+1})|\pi_{\text{new}}(\cdot|s_{t+1})\big{)}}\big{]}$ (20)
where $\mathcal{P}(s_{t},\mathbf{u_{t}},\cdot)$ is the transition function and $C=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ \Delta_{Q}(s_{t+1},\mathbf{u}_{t+1})=\underset{\mathbf{u}_{t+1}}{\mathrm{max}}\ |Q_{tot}^{\pi_{\text{old}}}(s_{t+1},\mathbf{u}_{t+1})-\sum_{i}\text{OPT}_{i}(Q_{i}^{\pi_{\text{old}}}(o^{i}_{t+1},u^{i}_{t+1}))|$.